A train travels 600 miles in the same time that a truck travels 520 miles. The train's average speed is 8 miles per hour faster than the truck's average speed. What is the speed of the train?
Answer by stanbon(75887)

A train travels 600 miles in the same time that a truck travels 520 miles.

The train's average speed is 8 miles per hour faster than the truck's average speed.

What is the speed of the train?
--------------------------
Train DATA:
distance = 600 miles ; rate = x+8 mph ; time = d/r = 600/(x+8) hrs.
--------------------------
Truck DATA:
distance = 520 miles ; rate = x mph ; time = d/r = 520/x hrs.
--------------------------
EQUATION:
time = time
600/(x+8) = 520/x
600x = 520(x+8)
15x = 13(x+8)
15x = 13x + 104
2x = 104
x = 52 mph (speed of the truck)
x+8 = 60 mph (speed of the train)
=====================
Cheers,
Stan H.